OpenStudy (anonymous):

Which of the following is the arithmetic sequence formula for this list? 16, 31, 46, 61, ... A. tn = 15n - 1 B. tn = 16n + 31 C. tn = n + 16

jimthompson5910 (jim_thompson5910):

What's the first term?

OpenStudy (anonymous):

Not sure :/

jimthompson5910 (jim_thompson5910):

It's just the left most term in that list

OpenStudy (anonymous):

16

jimthompson5910 (jim_thompson5910):

Good, now what's the common difference?

OpenStudy (anonymous):

I dont know :(

jimthompson5910 (jim_thompson5910):

Pick any term you want (except for the first one). Which term did you pick?

OpenStudy (anonymous):

31 i guess

jimthompson5910 (jim_thompson5910):

Now subtract off the previous term from it So 31 - 16 = ?

OpenStudy (anonymous):

15

jimthompson5910 (jim_thompson5910):

Good

jimthompson5910 (jim_thompson5910):

So the common difference is 15.

jimthompson5910 (jim_thompson5910):

The nth term is nth term = (first term) + (n-1)*(common difference) or tn = t1 + (n-1)*d where tn is the nth term, t1 is the first term, and d is the common difference

jimthompson5910 (jim_thompson5910):

From here, just plug in what you found tn = t1 + (n-1)*d tn = 16 + (n-1)*15 and simplify

OpenStudy (anonymous):

Well it say's 15m +1 But that's not an option

OpenStudy (anonymous):

n*

jimthompson5910 (jim_thompson5910):

hmm interesting, so there are only three options?

OpenStudy (anonymous):

Yes :/

jimthompson5910 (jim_thompson5910):

tn = 15n + 1 is the answer

OpenStudy (anonymous):

that's not an option, what do i do?

jimthompson5910 (jim_thompson5910):

ask your teacher about it, that's about all you can do really

jimthompson5910 (jim_thompson5910):

for now, just pick any other option (the options you listed are wrong, but oh well)

OpenStudy (anonymous):

I checked with my teacher and she said that the answer is listed

jimthompson5910 (jim_thompson5910):

well there must be a typo with the problem then

OpenStudy (anonymous):

What would be your best guess on the answer? This is my last test of the day and im trying to get it done so it can be SUMMER :D

jimthompson5910 (jim_thompson5910):

tn = 15n - 1 is the closest you get, but that -1 should be a +1

jimthompson5910 (jim_thompson5910):

So maybe the typo is there

OpenStudy (anonymous):

Alright, ill try that, thanks!

jimthompson5910 (jim_thompson5910):

sure thing